# DPLib: A Standard Benchmark Library for Distributed Power System Analysis and Optimization

### v2

@@ Abstract
DPLib is an open-source MATLAB-based benchmark library created to support research and development in distributed and decentralized power system analysis and optimization. Distributed and decentralized methods offer scalability, privacy preservation, and resilience to single points of failure, making them increasingly important for modern power systems. However, unlike centralized tools such as MATPOWER, no general-purpose, reproducible data library package currently exists for distributed power system studies. DPLib, available at GitHub, fills this gap by providing a standard power system library featuring over 20 multi-region benchmark test cases of varying sizes, along with a graph-based partitioning toolkit that decomposes any MATPOWER test system into multiple electrically coherent regions. The partitioning toolkit, an easy-to-use MATLAB code, generates standardized .mat and .m files, along with region visualizations for intuitive understanding. We also provide modular, easy-to-use distributed optimal power flow (OPF) solvers: an alternating direction method of multipliers(ADMM)-based DC-OPF solver implemented in YALMIP, and an ADMM-based AC-OPF solver leveraging IPOPT. These solvers validate the generated test systems for distributed optimization applications. Numerical results validate the generated test cases, establishing DPLib as a foundation for reproducible distributed power system research.
Index Terms:
Distributed power grid test systems, standard benchmark library, MATPOWER, distributed optimization.
@@ I Introduction
Centralized frameworks have long served as the dominant approach for modeling and solving large-scale power system problems, including optimal power flow (OPF) and security-constrained unit commitment. A key enabler behind the success and widespread adoption of centralized optimization and analysis in power systems is the availability of standardized benchmark test cases and libraries. Among these, MATPOWER has emerged as a widely accepted benchmark tool, offering consistent, well-structured data for centralized OPF studies  [1]. Its role in enabling reproducibility, fair performance comparisons, and methodological development has been fundamental to the evolution of centralized power system optimization and analysis.
For centralized OPF studies, for instance, [2] proposed convex relaxations for AC OPF and validated their results on MATPOWER test cases. [3] provided a comprehensive review of relaxation techniques for OPF, with all benchmarks sourced from MATPOWER. [4] introduced QC relaxation for OPF using MATPOWER data. More recently, [5] developed robust optimization methods for corrective control under uncertainty, relying on MATPOWER cases for validation. [6] proposed the NESTA archive to enrich MATPOWER with more realistic data, yet retained full MATPOWER compatibility. These works and many more underscore the importance of a common, reliable data standard—facilitated by MATPOWER—that allows researchers to validate and compare centralized optimization algorithms under consistent and reproducible conditions.
In the last decade, distributed and decentralized computation has gained significant momentum in the power systems community, especially in applications such as distributed OPF, voltage control, state estimation, and demand-side management [7, 8, 9]. These include studies that apply alternating direction method of multipliers (ADMM) decomposition [8, 10], distributed voltage control [11], parallel and distributed OPF [12, 13], consensus-based optimization [14], and cooperative control frameworks [15, 16]. Additional contributions include scalable distributed frameworks for transmission networks [17], graph-based coordination strategies [7], and distributed control in microgrids [18]. These methods offer inherent advantages in scalability, privacy preservation, and resilience to single points of failure.
@@
In response to this need, we develop a distributed power system library, DPLib, that addresses the lack of standardized resources for distributed power system studies by providing a comprehensive library of over 20 benchmark test cases, ranging from 73 to 9241 buses. These multi-region systems are generated using the DPLib graph-based partitioning toolkit that ensures electrical coherence and captures the corresponding multi-regional network topology. Each case includes the necessary structural and connectivity data to support distributed analysis. All test cases have been verified using distributed OPF problems to demonstrate their practicality and confirm their suitability for reproducible and scalable distributed optimization.
In addition to a standard library, a key requirement is a general and user-friendly partitioning tool capable of processing any standard MATPOWER case and dividing it into electrically coherent subregions based on user-defined parameters, such as the number of regions. An effective tool should preserve the integrity of the original network, identify and document inter-regional tie-lines, and produce output in a structured format consistent with centralized MATPOWER cases. This functionality enables users to generate customized test cases while maintaining compatibility with a standard data structure. The DPLib graph-based partitioning toolkit fulfills all of these requirements.
While the distributed library includes several pre-partitioned and validated datasets, researchers may also wish to generate and test their customized partitions using the DPLib graph-based toolkit. To support this, DPLib provides ADMM-based solvers for both DC OPF and AC OPF that are fully compatible with the partitioned data structure produced by the tool. These solvers take the partitioned datasets as input, iteratively solve the regional subproblems, enforce tie-line consistency through Lagrangian multipliers, and monitor convergence metrics, enabling reliable validation of user-defined partitions.
The main contribution of this work is the introduction of DPLib, a unified framework that establishes the first standard distributed data library for power system applications. Besides the library, DPLib consists of three core tools: a graph-based partitioning tool that decomposes any MATPOWER case into multiple regions using spectral clustering and k-means, enabling the creation of distributed datasets; a modular ADMM-based distributed DC OPF solver implemented in YALMIP; and a distributed AC OPF solver using ADMM and IPOPT. The DC OPF and AC OPF solvers are designed to validate and benchmark the partitioned datasets under realistic operating conditions. Together, these tools support reproducibility, scalability, and algorithm testing for distributed power system applications, addressing a key gap in current research infrastructure.
@@ II Graph-Based Partitioning and Toolkit Description
In many power systems studies, realistic standard test cases in the MATPOWER format are commonly used for testing and validation. This practice ensures that developed methods are evaluated on realistic cases that reflect the complexity of actual power systems. To maintain this standard, researchers working on distributed algorithms should also rely on realistic standard test systems and datasets rather than arbitrarily constructed or manually partitioned grids. In practice, real power networks often exhibit a naturally regional structure. For example, a network covering several regions will typically have strong internal connectivity within each region and only a few tie-lines connecting areas to one another. This results in regions that are strongly connected internally but weakly connected to the other areas. A proper partitioning method should be able to observe these loosely connected boundaries. We use graph-based partitioning to provide a principled and effective framework for capturing such structures and dividing the network in a way that aligns with real systems’ physical and topological characteristics.
@@ II-A Power Network Modeling and Spectral Clustering
To enable distributed optimization and control, a centralized network must first be divided into smaller regions that remain physically meaningful and electrically coherent. This partitioning process starts by modeling the power system as an undirected graph. We represent the power network as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where:
  • $\mathcal{V}$ is the set of buses (nodes), as defined in the original MATPOWER case file.
  • $\mathcal{E}$ is the set of transmission lines (edges), each connecting two buses.
The graph is undirected because the power flow direction does not impact the structure for partitioning purposes—only a connection between buses matters.
A path in a graph is a sequence of nodes such that an edge connects each consecutive pair of nodes in the sequence. A connected graph is an undirected graph with a path between every pair of nodes. This means it is possible to travel from any node to any other by following a sequence of edges in the graph. To describe the structure of $\mathcal{G}$, we construct the adjacency matrix $A\in\mathbb{R}^{n\times n}$, where $n=|\mathcal{V}|$ is the total number of buses. The entries of this matrix are defined as:
$A_{ij}=\begin{cases}1,&\text{if there is a line connecting bus }i\text{ and bus }j,\\ 0,&\text{otherwise}.\end{cases}$
Because the graph is undirected, $A$ is symmetric. Next, we define the degree matrix $D\in\mathbb{R}^{n\times n}$, a diagonal matrix where each diagonal entry $D_{ii}$ equals the number of lines connected to bus $i$:
$D_{ii}=\sum_{j=1}^{n}A_{ij}.$
Using these two matrices, we define the graph Laplacian $L=D-A.$ The Laplacian matrix $L$ encodes the power network connectivity and topology. For a connected graph, $L$ is symmetric and positive semi-definite, meaning all its eigenvalues are non-negative. Importantly, the smallest eigenvalue of $L$ is always zero, corresponding to the trivial uniform eigenvector.
To divide the system into $k$ electrically coherent regions, we use spectral clustering, which leverages the properties of $L$. The first step is to compute the eigenvalues and eigenvectors of $L$. Among these, the $k$ smallest nonzero eigenvalues are selected. Each eigenvalue is associated with an eigenvector that varies smoothly across the graph, meaning that buses that are strongly connected in the physical system will have similar eigenvalues. Therefore, these eigenvectors reflect the underlying structure of the network.
@@
The clustering result yields a partition of all buses into $k$ regions. Each cluster groups strongly connected buses, leading to electrically coherent and balanced regions. After forming the clusters, inter-regional tie-lines (lines that connect buses belonging to different clusters) are identified. To improve quality, the algorithm may repeat the clustering multiple times and select the configuration that minimizes the number of inter-regional tie-lines.
This spectral clustering approach provides a systematic and scalable method for decomposing any MATPOWER case into meaningful regions. It aligns well with practical expectations since most real-world partitioned power networks are formed based on geographic or electrical proximity.
A tool that can automatically convert MATPOWER test cases into $k$ well-structured regions is essential for distributed power system studies. The DPLib partitioning toolkit fills this gap by generating standardized multi-region datasets from any MATPOWER case, supporting reproducibility and scalability.
@@ II-B DPLib Partitioning Toolkit
The DPLib graph-based partitioning toolkit implements the method described above and automates the generation of partitioned MATPOWER cases. The partitioning toolkit is publicly available at GitHub, including the MATLAB implementation file named partitioning_code.m. The code uses MATLAB R2020a or newer, along with the MATPOWER toolbox. Upon running the script, the user is prompted to enter the name of a valid MATPOWER case file (for example, case118) and specify the desired number of regions into which the system should be divided.
Internally, the code loads the MATPOWER case and checks whether the bus numbers are sequential. If not, it renumbers the buses from one to $n$ and updates all corresponding references in the branch and generator matrices. It then constructs the undirected adjacency matrix based on the network’s branch connectivity and forms the corresponding graph Laplacian. The eigenvectors of the Laplacian are computed and used as features for $k$-means clustering. The clustering process is repeated multiple times to ensure optimality, and the configuration with the lowest number of inter-regional tie-lines is selected.
For each candidate partitioning, regional subcases are formed by extracting internal buses, branches, generators, and generation cost data for each region. The script ensures that each region has at least one generator. Buses are then renumbered locally within each region to maintain consistency. Inter-regional tie-lines are identified as any branch connecting buses in different regions and stored.
Once a feasible partitioning is found, the code saves the result in multiple formats. A MATLAB .mat file is created, containing all regional subcases under separate variables (e.g., mpc_regionR1, mpc_regionR2, etc.), along with a comprehensive table of all inter-regional tie-lines labeled by their originating and receiving regions and complete branch data. A corresponding .m file is also generated, exporting the same structures in a human-readable script format suitable for inspection or direct use in other codes. Additionally, the code creates a graphical visualization of the partitioned system, where each region is displayed as a labeled node in a circular layout, and tie-lines are represented as edges connecting the areas. The number of tie-lines between each pair of regions is also annotated.
During execution, the code prints detailed messages to the MATLAB command window. These include whether bus renumbering was necessary, confirmation of full graph connectivity, the number of generated regions and tie-lines, the region assignment of the reference (slack) bus, and whether all buses, branches, and generators are fully included in the partitioned result. It also confirms the successful saving of the generated .mat, .m, and .png files. The algorithm of the partitioning toolkit, illustrating the complete data processing pipeline, is provided in 1.
@@
1:User Input:
  • Name of the original MATPOWER case (e.g., case118)
  • Number of desired regions $k$
2: Load the MATPOWER case file
3: Check bus numbering; if non-sequential, renumber from 1 to $n$ and update all references
4: Construct adjacency matrix $A$ from the branch data.
5: Construct degree matrix $D$ with $D_{ii}=\sum_{j}A_{ij}$
6: Form graph Laplacian $L=D-A$
7: Compute the first $k$ nonzero eigenvectors of $L$
8: Stack eigenvectors into matrix $U\in\mathbb{R}^{n\times k}$
9: Apply $k$-means clustering to the rows of $U$
10: Repeat clustering for multiple iterations to minimize the number of tie-lines
11:for each region (cluster) do
12:  Extract internal buses, branches, generators, and cost data
13:  Ensure that at least one generator exists in each region
14:  Locally renumber buses
15:  Identify inter-regional tie-lines
16:end for
17: Save output in the following formats:
  • .mat file with region-wise MATPOWER data
  • .m script file exporting the same data
  • .png visualization showing regions and tie-lines
@@ III Validating DPLib Test Systems Using OPF
To ensure that the generated multi-region test cases are electrically coherent and suitable for distributed power system studies, it is essential to validate them through power system analysis. Although the DPLib partitioning toolkit generates distributed variants of MATPOWER test systems applicable to a wide range of power system studies, including control, state estimation, or stability analysis, this section evaluates these test systems for distributed OPF applications. OPF has been selected as a representative benchmarking problem because of its maturity, availability of centralized references, and compatibility with distributed solvers due to over a decade of research in this area. We have developed general ADMM-based solvers for DC and AC OPF problems that operate directly on the test cases produced by the DPLib partitioning toolkit. These solvers are designed to validate the feasibility, consistency, and convergence of partitioned datasets, demonstrating the suitability of these test systems. Nevertheless, users may use alternative solvers or utilize these provided tools to verify and explore their own partitioned datasets.

### v3

@@ Abstract
DPLib is an open-source MATLAB-based benchmark library created to support research and development in distributed and decentralized power system analysis and optimization. Distributed and decentralized methods offer scalability, privacy preservation, and resilience to single points of failure, making them increasingly important for modern power systems. However, unlike centralized tools such as PGLib, no general-purpose, reproducible data library package currently exists for distributed power system studies. DPLib, available at GitHub, fills this gap by providing a standard power system library featuring over 20 multi-region benchmark test cases of varying sizes, along with a graph-based partitioning toolkit that decomposes any MATPOWER test system into multiple regions. The partitioning toolkit, an easy-to-use MATLAB code, generates standardized .mat and .m files, along with region visualizations for intuitive understanding. We also provide modular, easy-to-use distributed optimal power flow (OPF) solvers: an alternating direction method of multipliers (ADMM)-based DC OPF solver implemented in YALMIP, and an ADMM-based AC OPF solver leveraging IPOPT. These solvers validate the generated test systems for distributed optimization applications. Numerical results validate the generated test cases, establishing DPLib as a foundation for reproducible distributed power system research.
@@ I Introduction
Centralized frameworks have long served as the dominant approach for modeling and solving large-scale power system problems, including optimal power flow (OPF), planning, control, and unit commitment. A key enabler behind the success and widespread adoption of centralized optimization and analysis in power systems is the availability of standardized benchmark test cases and libraries. Among these, MATPOWER has emerged as a widely accepted benchmark tool, offering consistent, well-structured data for centralized studies [22]. Its role in enabling reproducibility, fair performance comparisons, and methodological development has been fundamental to the evolution of centralized power system analysis.
For centralized studies, for instance, [11] proposed convex relaxations for AC OPF and validated their results on MATPOWER test cases. [16] provided a comprehensive review of relaxation techniques for OPF, with all benchmarks sourced from MATPOWER. [2] introduced QC relaxation for OPF using MATPOWER data. More recently, [19] developed robust optimization methods for corrective control under uncertainty, relying on MATPOWER cases for validation. [1] proposed the NESTA archive to enrich MATPOWER with more realistic data, yet retained full MATPOWER compatibility. These works and many more underscore the importance of a common, reliable data standard—facilitated by MATPOWER—that allows researchers to validate and compare centralized optimization algorithms under consistent and reproducible conditions.
In the last decade, distributed and decentralized computation has gained significant momentum in the power systems community, especially in applications such as distributed OPF, voltage control, state estimation, and demand-side management [15, 10, 6]. These include studies that apply alternating direction method of multipliers (ADMM) decomposition [10, 4], distributed voltage control [14], parallel and distributed OPF [21, 8], consensus-based optimization [18], and cooperative control frameworks [3, 7, 9]. Additional contributions include scalable distributed frameworks for transmission networks [12], graph-based coordination strategies [15], and distributed control in microgrids [5]. These methods offer inherent advantages in scalability, privacy preservation, and resilience to single points of failure.
@@
In response to this need, we develop a distributed power system library, DPLib, that addresses the lack of standardized resources for distributed power system studies by providing a comprehensive library of over 20 benchmark test cases, ranging from 73 to 9241 buses. These multi-region systems are generated using the DPLib graph-based partitioning toolkit that captures the corresponding multi-regional network topology. Each case includes the necessary structural and connectivity data to support distributed analysis. All test cases have been verified using distributed OPF problems to demonstrate their practicality and confirm their suitability for reproducible and scalable distributed optimization.
In addition to a standard library, a key requirement is a general and user-friendly partitioning tool capable of processing any standard MATPOWER case and dividing it into subregions based on user-defined parameters, such as the number of regions. An effective tool should preserve the integrity of the original network, identify and document inter-regional tie-lines, and produce output in a structured format consistent with centralized MATPOWER cases. This functionality enables users to generate customized test cases while maintaining compatibility with a standard data structure. The DPLib graph-based partitioning toolkit fulfills all of these requirements.
While the distributed library includes several pre-partitioned and validated datasets, researchers may also wish to test their customized partitions generated by the DPLib graph-based toolkit. To support this, DPLib provides ADMM-based solvers for both DC OPF and AC OPF that are fully compatible with the partitioned data structure produced by the tool. These solvers take the partitioned datasets as input, iteratively solve the regional subproblems, enforce tie-line consistency through Lagrangian multipliers, and monitor convergence metrics, enabling reliable validation of user-defined partitions.
The main contribution of this work is the introduction of DPLib, a unified framework that establishes the first standard distributed data library for power system applications. Besides the library, DPLib consists of three core tools: a graph-based partitioning tool that decomposes any MATPOWER case into multiple regions using spectral clustering and k-means, enabling the creation of distributed datasets; a modular ADMM-based distributed DC OPF solver implemented in YALMIP [13]; and a distributed AC OPF solver using ADMM and IPOPT [20]. The DC OPF and AC OPF solvers are designed to validate and benchmark the partitioned datasets under realistic operating conditions. Together, these tools support reproducibility, scalability, and algorithm testing for distributed power system applications, addressing a key gap in current research infrastructure.
@@ II Graph-Based Partitioning and Toolkit Description
In many power systems studies, realistic standard test cases in the MATPOWER format are commonly used for testing and validation. This practice ensures that developed methods are evaluated on realistic cases that reflect the complexity of actual power systems. To maintain this standard, researchers working on distributed algorithms should also rely on realistic standard test systems and datasets rather than arbitrarily constructed or manually partitioned grids.
In practice, real power networks often exhibit a naturally regional structure. For example, a network covering several regions will typically have strong internal connectivity within each region and only a few tie-lines connecting areas to one another. This results in regions that are strongly connected internally but weakly connected to the other areas. A proper partitioning method should be able to observe these loosely connected boundaries. We use graph-based partitioning to provide a principled and effective framework for capturing such structures and dividing the network in a way that aligns with real systems’ physical and topological characteristics.
@@ II-A Power Network Modeling and Spectral Clustering
To prepare the network for spectral partitioning, the power system is modeled as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where buses form the nodes and transmission lines form the edges. The graph representation is encoded through the adjacency matrix $A$, and two forms of this matrix may be constructed depending on how connectivity is interpreted.
In its simplest form, the network is viewed purely through its topological structure. Each transmission line indicates the presence of a connection, and the adjacency matrix is as
$A_{ij}=\begin{cases}1,&\text{if a line connects buses }i\text{ and }j,\\ 0,&\text{otherwise}.\end{cases}$ (1)
This unweighted representation treats all lines uniformly and emphasizes the combinatorial shape of the network. Such a formulation is robust to data irregularities, reflects the way regional boundaries appear in practice when tie-lines are sparse, and preserves the graph-theoretic properties that spectral clustering exploits most directly.
A second option introduces electrical heterogeneity explicitly. In this case, the weight of an edge is determined by parameters such as line reactance $x_{ij}$, line charging susceptance $b_{ij}$, and thermal rating $S_{\max,ij}$. After appropriate normalization, a composite quantity is assigned to each existing edge,
$A_{ij}=w(x_{ij},b_{ij},S_{\max,ij}),$ (2)
so the graph captures not only the presence of a connection but also its relative electrical strength. This weighted representation highlights tightly coupled corridors and down-weights weak or limited-capacity lines, and therefore encodes additional physical detail into the graph used for clustering.
Although the toolkit supports a weighted Laplacian that incorporates electrical heterogeneity, we treat this option as an auxiliary feature rather than the default mode. Weighted formulations can highlight electrically strong corridors, but doing so requires a choice of normalization, scaling, and aggregation of heterogeneous parameters—decisions that are not standardized across power-system studies. Because the purpose of DPLib is to produce baseline multi-region datasets for benchmarking distributed algorithms, we use the unweighted formulation to ensure that all results are comparable, repeatable, and free of case-dependent parameter tuning. Researchers who wish to emphasize electrical strength rather than topological connectivity may enable the weighted variant, which the toolkit provides for completeness.
Both constructions lead to the degree matrix $D$ defined by $D_{ii}=\sum_{j}A_{ij}$, and the Laplacian L=DA.L=D-A. The Laplacian matrix $L$ encodes the power network connectivity and topology. For a connected graph, $L$ is symmetric and positive semi-definite, meaning all its eigenvalues are non-negative. Importantly, the smallest eigenvalue of $L$ is always zero, corresponding to the trivial uniform eigenvector.
To divide the system into $k$ regions, we use spectral clustering, which leverages the properties of $L$. The first step is to compute the eigenvalues and eigenvectors of $L$. Among these, the $k$ smallest nonzero eigenvalues are selected. Each eigenvalue is associated with an eigenvector that varies smoothly across the graph, meaning that buses that are strongly connected in the physical system will have similar eigenvalues. Therefore, these eigenvectors reflect the underlying structure of the network.
@@
The clustering result yields a partition of all buses into $k$ regions. Each cluster groups strongly connected buses, leading to balanced regions where each has its generation. After forming the clusters, inter-regional tie-lines (lines that connect buses belonging to different clusters) are identified. To improve quality, the algorithm repeats the clustering multiple times and select the configuration that minimizes the number of inter-regional tie-lines.
@@ II-B DPLib Partitioning Toolkit
The DPLib partitioning toolkit implements the graph-based clustering framework described in the previous subsection and automates the full workflow for constructing multi-region MATPOWER datasets. The toolkit supports both adjacency representations introduced earlier, allowing the Laplacian to be formed either from the purely topological structure of the network or from a weighted graph that incorporates electrical parameters. In the unweighted mode, each transmission line contributes a unit entry to the adjacency matrix, so the Laplacian reflects the combinatorial connectivity of the system. In the weighted mode, each existing edge is assigned a composite weight constructed from line reactance, line charging susceptance, and thermal capacity after normalization, allowing the Laplacian to embed relative electrical strength directly into the spectral geometry.
Although both modes are supported, the unweighted Laplacian is used as the default for generating standardized multi-region test cases. This choice ensures that the resulting partitions depend only on the physical topology of the MATPOWER system and remain independent of parameter scaling, model variations, or dataset-dependent numerical ranges. The unweighted construction also aligns with how regional boundaries appear in real-world transmission networks, where geographical separations and sparse tie-lines primarily determine structural decomposition. Using the unweighted version therefore provides consistent, reproducible partitions across different datasets and preserves compatibility with studies that compare methods on topologically equivalent cases.
The workflow proceeds as follows. The MATPOWER case is loaded, and if the original bus numbering is not sequential, all buses, branches, and generators are renumbered from $1$ to $n$. Depending on the selected mode, the adjacency matrix is constructed using either the binary definition or the weighted formulation. The degree matrix and Laplacian are then formed, and the $k$ Laplacian eigenvectors associated with the smallest nonzero eigenvalues are computed to obtain the spectral embedding. Multiple runs of $k$-means are applied to the rows of this matrix to obtain a stable clustering, and the configuration with the smallest number of inter-regional tie-lines is selected. For the chosen partition, regional subcases are created by extracting internal buses, branches, generators, and cost data, and each region is locally renumbered to maintain consistency. Tie-lines connecting different regions are identified and stored together with their full MATPOWER branch data.
The completed partition is exported in several formats. A .mat file contains all regional subcases and a table describing every inter-regional tie-line. A corresponding .m script reproduces the same structures in a clear and readable format. A graphical visualization is produced in .png format, showing regions arranged in a force-directed layout together with annotated tie-lines. These outputs provide a consistent, reproducible, and easily verifiable multi-region dataset that can be used directly in distributed studies. The algorithm of the partitioning toolkit, illustrating the complete data processing pipeline, is provided in Algorithm 1.
@@
1: Input MATPOWER case name and desired number of regions $k$
2: Load MATPOWER data; if bus numbering is non-sequential, renumber buses and update references
3: Construct adjacency matrix AA:
  1. 1.
    Unweighted: $A_{ij}=1$ if a branch connects $i$ and $j$
  2. 2.
    Weighted: assign $A_{ij}=w(x_{ij},b_{ij},S_{\max,ij})$
4: Form degree matrix $D$ and Laplacian L=DAL=D-A
5: Compute the Laplacian eigenvectors associated with the $k$ smallest nonzero eigenvalues and form $U$
6: For multiple attempts: apply $k$-means to rows of $U$, build candidate regions, identify tie-lines, and discard candidates without generators
7: Select the valid partition with the minimum number of tie-lines
8:for each region do
9:  Extract internal buses, branches, generators, and cost data; locally renumber buses
10:end for
11: Identify the region of the slack bus and record all inter-regional tie-lines
12: Export regional cases to .mat and .m files and generate topology visualization
@@ III Validating DPLib Test Systems Using OPF
To ensure that the generated multi-region test cases are suitable for distributed power system studies, it is essential to validate them through power system analysis. Although the DPLib partitioning toolkit generates distributed variants of MATPOWER test systems applicable to a wide range of power system studies, including control, state estimation, or stability analysis, this section evaluates these test systems for distributed OPF applications only. OPF has been selected as a representative benchmarking problem because of its maturity, availability of centralized references, and compatibility with distributed solvers due to over a decade of research in this area. We have developed general ADMM-based solvers for DC and AC OPF problems that operate directly on the test cases produced by the DPLib partitioning toolkit. These solvers are designed to validate the feasibility, consistency, and convergence of partitioned datasets, demonstrating the suitability of these test systems. Nevertheless, users may use alternative solvers or utilize these provided tools to verify and explore their own partitioned datasets.